# Towards Data-Efficient Pretraining for Atomic Property Prediction

### v1

@@ Appendix A More Epochs or More Data?
Appendix B Additional Analysis on the Metric Design
Figure 6: Impact of using mean aggregation instead of flattening on CSI values. We notice that the mean pooling incorrectly reduced the score for OC22 potentially due to over-smoothing.
Figure 7: Impact of using random sampling strategy instead of class-balanced sampling. As highlighted in the long-tail analysis in Appendix C, random sampling can lead to class underrepresentation, potentially affecting the correlation between upstream and downstream tasks. Notably, both ANI-1x and Transition-1x exhibit different patterns compared to the class-balanced values reported in the main paper.
Figure 8: The impact of using another backbone. We use JMP pretrained model and show that similar insights are obtained where Ani-1x is shown as the most similar.
Appendix C Long-Tail Analysis
As discussed in Section 3, molecular understanding datasets are typically imbalanced and exhibit a skewed distribution. Figure 9 illustrates the effect of random sampling versus class-balanced sampling, which we employ in this study, in terms of class coverage. Notably, random sampling results in significant class underrepresentation, whereas class-balanced sampling ensures broader coverage across all classes.
@@ Appendix C Long-Tail Analysis
Figure 9: Impact of sampling strategies on subset construction for feature extraction. We sample 10K instances for each upstream task, highlighting the differences in class coverage between random and class-balanced sampling.
Appendix D Implementation Details

### v2

@@ Appendix A More Epochs or More Data?
Appendix B Impact of Changing the Feature Extractor for CSI
In the main paper, we used an EquiformerV2 model pretrained on the OC20 dataset as our primary feature extractor for computing the CSI values. In this section, we explore alternative open-source pretrained EquiformerV2 checkpoints, including: ODAC23 (Sriram et al., 2024), MPtrj (Barroso-Luque et al., 2024) and MPtrj (DeNS) (Barroso-Luque et al., 2024). DeNS (Liao et al., 2024) refers to a denoising auxiliary task that extends traditional denoising to non-equilibrium structures to improve learned representations.
We report the results in Figure 7. Our comparison is based on relative rankings since the absolute values are influenced by the choice of feature extractor. Notably, for 4 out of 5 datasets (rMD17, MD22, SPICE, and OMat24), all feature extractors agree on the lowest-CSI upstream dataset: ANI-1x for rMD17, MD22, and SPICE, and OC20 for OMat24. For QM9, OC20 and MPtrj (DeNS) rank ANI-1x as the most aligned, whereas ODAC23 and MPtrj place ANI-1x second to Transition1x, which is also a strong upstream candidate for QM9.
Figure 7: Dataset Alignment Using Feature Extractors Pretrained on Various Datasets. We compare CSI relative values computed using feature extractors pretrained on four different datasets: OC20, ODAC23, MPtrj, and MPtrj (DeNS). The lowest CSI upstream dataset is consistent across feature extractors for 4 out of 5 downstream tasks, namely rMD17, MD22, SPICE, and OMat24.
Appendix C Feature Representation Choices for CSI
A key challenge when adapting the Fréchet Inception Distance (FID) to 3D molecular graphs is handling their variable number of atoms. Our feature extractor, EquiformerV2, produces variable-sized node-level features in $\mathbb{R}^{n\times d}$, where $n$ is the number of atoms and $d$ is the feature dimension. A straightforward approach is to aggregate these features using mean pooling, producing a graph-level representation in $\mathbb{R}^{N\times d}$, where $N$ is the number of graphs. This setup is directly analogous to the image-based FID. The central question, however, is whether this aggregation loses valuable information.
For our CSI metric to accurately measure alignment across various domains, it is critical that the feature expressivity is maximized. To quantify this expressivity, we use the Effective Rank (Roy and Vetterli, 2007) of the covariance matrix of these features. The Effective Rank is similar to the traditional matrix rank but is more robust to noise.
Formally, as in  Roy and Vetterli (2007), let $C\in\mathbb{R}^{d\times d}$ be the covariance matrix with eigenvalues $\{\lambda_{j}\}_{j=1}^{d}$ and $p_{j}=\lambda_{j}/\sum_{m=1}^{d}\lambda_{m}$. The Shannon (1948) entropy is $H=-\sum_{j=1}^{d}p_{j}\log p_{j}$, and the Effective Rank is:
$\operatorname{erank}(C)=\exp(H).$
Higher $\operatorname{erank}$ indicates that information is distributed across more independent directions, signifying a richer and less redundant feature representation.
To compare node-level and graph-level expressivity in a paired manner, we employ a bootstrap protocol. For a chosen number of graphs $k$, we proceed as follows:
  1. 1.
    Sample $k$ graphs from the dataset.
  2. 2.
    Construct two matrices in $\mathbb{R}^{k\times d}$ from the sampled graphs: (i) a graph-level matrix, where each row is the pooled representation of one graph, and (ii) a node-level matrix, where each row corresponds to a randomly selected node representation from each graph.
  3. 3.
    Standardize each feature (z-score per dimension) in both matrices.
  4. 4.
    Compute the covariance $C$ and its effective rank $\operatorname{erank}(C)$ for both variants.
We repeat this process 10 times and report the mean effective ranks. We perform this comparison for $k\in\{5{,}000,10{,}000,15{,}000\}$ and show the results in Figure 8. We observe that across all tested sample sizes, the node-level features consistently have a significantly larger effective rank. This finding indicates that the aggregation process for graph-level features results in a loss of intrinsic dimensions that are likely significant for dataset alignment studies. Therefore, to preserve maximum feature expressivity, we use node-level features in our CSI design.
Figure 8: Node-Level vs Graph-Level Effective Rank. Effective rank of features from (ANI-1x, Transition-1x, OC20, OC22) datasets using both node-level (green) and mean-pooled graph-level (orange) representations. The consistently higher effective rank of node-level features demonstrates that mean pooling removes informative directions, which may be inadequate for analyzing alignment between datasets.
To complement our effective rank analysis, we compare node-level and graph-level CSI variants in Figure 9. OC20 and OC22 are both catalysis-focused datasets and share similar chemical domains. The node-level features reflect this similarity by yielding comparable scores between the two datasets. In contrast, the graph-level variant suggests greater divergence between the two, which may be due to information loss introduced by mean pooling.
Figure 9: CSI values for Node-Level vs Graph-Level Feature Representation
Appendix D Class Balancing for Feature Extraction
The upstream datasets analyzed in our main paper are large in size. For example, OC20 contains over 133 million training samples, while Transition-1x contains around 9.6 million. To make the computation of our CSI metric tractable, we sample 10K instances from each upstream dataset for feature extraction. To ensure these subsets are representative of the full distribution, we perform class-balanced sampling. For ANI-1x and Transition-1x, a class is defined by the molecular formula of the molecule in the trajectory. For OC20 and OC22, a class is defined by the bulk structure. Each class is treated as a bin, and we sample an equal number of instances from each bin.
Figure 10 compares random sampling and class-balanced sampling in terms of class coverage. Random sampling leads to significant underrepresentation of less frequent classes, while class-balanced sampling ensures uniform coverage across all classes.
@@ Appendix D Class Balancing for Feature Extraction
Figure 10: Impact of Sampling Strategies on Subset Construction for Feature Extraction. We show the differences in class coverage between random and class-balanced sampling when sampling 10K instances for each upstream task.
Appendix E Implementation Details